OpenStudy (anonymous):

Use the following equation: log2x + log2(x – 7) = 3 The third step would be to: Select one: a. Graph the function. b. Take the log of each side. c. Square the expression. d. Factor the resulting quadratic equation.

jhonyy9 (jhonyy9):

do you know the properties of logarithm ?

jhonyy9 (jhonyy9):

so log a +log b = log (a*b)

jhonyy9 (jhonyy9):

so and 3 how you can rewrite it using log2 ?

jhonyy9 (jhonyy9):

3=log2 2^3 =log2 8

jhonyy9 (jhonyy9):

so than log2 x +log2 (x-7) =3 log2 (x(x-7)) = log2 8 ok ?

jhonyy9 (jhonyy9):

so x(x-7)=8 x^2 -7x -8 =0